f1= 97base10

f2= -326base10

A)add f1 to f2 and convert the solution to octal, hex and binary

B)what is the range and quantity of a 12 bit signed register
Answer #1

(A)f1 + f2 = (97 + (-326)) = -229
Binary (2's complement) = 100011011
octal = 433
hexadecimal = 0xF1B

(B)Range of 1 bit signed number
-(2^11) to (2^11 - 1)
=> -2048 to 2047
quantity = 2^12 = 4096